# Scalable and Cost-Efficient de Novo Template-Based Molecular Generation

### v1

@@ 4.6 Synthesibility-Oriented Metrics
We compare SCENT with leading GFlowNet baselines for molecular generation: RGFN and the fragment-based GFlowNet (FGFN) from Bengio et al. [7]. We also evaluate FGFN (SA), trained to jointly optimize the sEH proxy and the Synthetic Accessibility (SA) score. As shown in Figure 7, template-based models substantially improve synthesizability. This is most evident in the AiZynthFinder success rate [48], where FGFN fails completely and FGFN (SA) performs poorly relative to RGFN and SCENT. Notably, although FGFN (SA) achieves the lowest average SA score, it still struggles on AiZynthFinder, underscoring the limitations of heuristic synthesizability metrics. SCENT achieves the highest average reward and outperforms RGFN in synthesizability, which we attribute to its preference for shorter reaction pathways (Section 4.3), resulting in reduced molecular weight that correlates with the synthesibility metrics [49].
@@ 5 Conclusion
In this paper, we presented SCENT, a template-based approach for molecular generation within the GFlowNet framework. SCENT introduces three new mechanisms that collectively enhance exploration of chemical space: (1) Resursive Cost Guidance, a general strategy for guiding GFlowNet’s backward policy, which we utilized to improve computational efficiency by removing costly parent state computation, and to focus sampling on molecules with low synthesis cost, (2) Dynamic Library building blocks, which improve exploration efficiency—especially with smaller fragment libraries—and enable the generation of non-linear synthesis trees, and (3) an Exploration Penalty, which avoids repeated visits to previously explored candidates. Through empirical evaluation, we demonstrated that combining these strategies leads to state-of-the-art results in template-based molecular generation across diverse experimental settings, from small-scale fragment libraries suitable for rapid, automated synthesis to large-scale libraries that create vast search spaces.
@@ A.3.1 Synthesis Cost Model
We train the model $\hat{c}_{B}^{S}$ using the procedure described in Section 3.2. Given an updated dataset with synthesis cost estimates $D$, we perform five mini-batch updates of $\hat{c}_{B}^{S}$ using mini-batches of size $1024$, the learning rate of $0.01$ and the mean square error loss $\mathcal{L}_{cost}=\sum_{c_{s}\in D^{\prime}}\lvert\lvert c_{s}-\hat{c}_{B}^{S}\rvert\rvert_{2}^{2}$, where $D^{\prime}$ is a mini-batch of $D$. To stabilize training, we standardize all costs in $D$ using the mean and variance of costs from the building block library $M$.
@@ A.3.2 Decomposability Model
The decomposability model $\hat{c}_{B}^{D}$ is very similar to the synthesis cost model, except that it encodes the current number of reactions used in the molecule $s$. The "cost" in the context of decomposability is 00 if the given molecule can be decomposed into building blocks from $M$ (within some number of steps) and $1$ otherwise. In practice, we train the model $\hat{c}_{B}^{D}$ to predict the decomposability label $\in\{0,1\}$ using binary cross entropy $\mathcal{L}_{cost}$. The labels are gathered similarly to the synthesis cost values, using the trajectories encountered in the GFlowNet forward sampling. We update $\hat{d}(s)$ with five minibatches of size $1024$ and a learning rate of $0.005$. In addition, we ensure that at least $20\%$ of the samples of molecules for training $\hat{d}(s)$ are not decomposable.
@@ B.1 SCENT
All SCENT instances were trained using Decomposability Guidance, either solely or with Synthesis Cost Guidance using Equation 11.
@@
To predict the decomposability of the molecule, we use the model described in Section A.3. We set the cost temperature (Equation 10) to $\alpha_{D}=5$.
@@
The model for Synthesis Cost Guidance is described in Section A.3. We set the cost temperature to $\alpha=5$.
@@
We set $\epsilon=3$ in Equation 12 and schedule $\gamma$ in two ways: 1) so that it linearly decays to zero after $N$ iterations, and 2) it grows with the trajectory length (each step increases it by some constant factor $\Delta\gamma$). We set $N=\infty$ for SCENT that uses Dynamic Library and $N=1000$ for the rest of the runs. The growing delta $\Delta\gamma=0.2$, while the initial $s_{0}$ temperature $\gamma_{0}=1.0$. Note that for JNK3 in the MEDIUM setting, we used $4500$ iterations for exploitation penalty.
@@ B.4 RxnFlow
Appendix C Reaction Yields and Fragments Costs Estimates
@@
Building block prices for SMALL were obtained from the vendors online. If multiple vendors offered the same building block, only the cost of the cheapest option was considered. For compounds available in varying amounts, both the price per gram and the alternative size (e.g., 5, 10, 25 grams) were considered, and the smallest price per mmol was used. Compounds that were exclusively accessible in other forms, such as salts, related functional groups, or containing a protected functional group were used with the SMILES fitted to the corresponding reactions. The cost of a product includes the total cost of the building blocks used, while additional expenses such as solvents, catalysts, and reagents are not accounted for.
Reaction yield estimates for amide coupling, nucleophilic additions to isocyanates, Suzuki reaction, Buchwald-Hartwig coupling, Sonogashira coupling, and azide-alkyne cycloaddition were obtained by analyzing in-house data. For all other reactions, SMARTS templates were used to search for reaction substructure in Reaxys and SciFinder, and the obtained literature data was analyzed to produce yield estimates.
@@
Appendix D Extended Comparison to Template-Based GFlowNets
D.1 Results on sEH with Reward Threshold = 7.2
Despite tuning the SynFlowNet [9] parameters (Appendix B), we could not obtain satisfying results. Importantly, our experimental setting differs from the one used in their paper in the following way: 1) we use up to 4 reactions, they use 2 or 3, 2) they use a 10k subset of Enamine, our MEDIUM setting contains 64k molecules, 3) we set the reward thresholds higher (for sEH we set 8.0 instead of 7.2), and 4) we train the model for fewer trajectories. In Table 2, we report the results for sEH using a threshold from the SynFlowNet paper. The number of modes obtained seems consistent with those reported in Cretu et al. [9].
Table 2: Evaluation of different template-based approaches using sEH proxy in SMALL, MEDIUM and LARGE settings, but using lower reward threshold ($>7.2$) equivalent to the one used in SynFlowNet paper [9].
@@ Appendix D Extended Comparison to Template-Based GFlowNets
modes $>$ 7.2 $\uparrow$ scaff. $>$ 7.2 $\uparrow$ reward $\uparrow$ cost $>$ 7.2 $\downarrow$
SMALL RGFN 11320 $\pm$168 76742 $\pm$693 7.34 $\pm$0.06 35.3 $\pm$1.6
@@
modes $>$ 0.8 $\uparrow$ scaff. $>$ 0.8 $\uparrow$ reward $\uparrow$ cost $>$ 0.8 $\downarrow$
SMALL RGFN 285 $\pm$14 1870 $\pm$106 0.7 $\pm$0.01 42.3 $\pm$2.7
@@
Table 4: Evaluation of different template-based approaches using JNK3 proxy in SMALL, MEDIUM and LARGE settings. In all metrics, except "reward", we consider molecules with reward $>0.8$. We observe that SCENT (C) that uses Cost Guidance can greatly reduce the cost of the sampled molecules in SMALL and MEDIUM settings, while drastically increasing the number of scaffolds.
@@
modes $>$ 0.8 $\uparrow$ scaff. $>$ 0.8 $\uparrow$ reward $\uparrow$ cost $>$ 0.8 $\downarrow$
SMALL RGFN 307 $\pm$11 2852 $\pm$42 0.8 $\pm$0.0 11.6 $\pm$0.3
@@ D.3 Results on JNK3
In Table 4 RGFN outperforms all SCENT variants in the smallest setting, consistently finding more diverse modes and high-reward scaffolds. While SCENT lags behind in terms of mode discovery, it compensates in terms of cost, since SCENT (C) samples the cheapest molecules. In the MEDIUM setting, SCENT (w/o C) outperforms the baseline methods and the fully enabled SCENT variants in terms of mode discovery, but lags behind in the number of high-reward scaffolds. That can be attributed to the fact that SCENT (C) samples are less diverse. However, SCENT (C+P) samples the cheapest molecules. Finally, in the largest setting, the fully enabled SCENT (C+P) finds the most number of modes and high-reward scaffolds, whereas SCENT (C) takes the spotlight in terms of cost, generating the cheapest molecules.
Appendix E Extended Synthesis Cost Guidance Analysis

### v2

@@ 4.6 Synthesibility-Oriented Metrics
We compare SCENT with leading GFlowNet baselines for molecular generation: RGFN and the fragment-based GFlowNet (FGFN) from Bengio et al. [7]. We also evaluate FGFN (SA), trained to jointly optimize the sEH proxy and the Synthetic Accessibility (SA) score. As shown in Figure˜7, template-based models substantially improve synthesizability. This is most evident in the AiZynthFinder success rate [48], where FGFN fails completely and FGFN (SA) performs poorly relative to RGFN and SCENT. Notably, although FGFN (SA) achieves the lowest average SA score, it still struggles on AiZynthFinder, underscoring the limitations of heuristic synthesizability metrics. SCENT achieves the highest average reward and outperforms RGFN in synthesizability, which we attribute to its preference for shorter reaction pathways (Section˜4.3), resulting in reduced molecular weight that correlates with the synthesibility metrics [49].
@@ 5 Conclusion
In this paper, we presented SCENT, a template-based approach for molecular generation within the GFlowNet framework. SCENT introduces three new mechanisms that collectively enhance exploration of chemical space: (1) Resursive Cost Guidance, a general strategy for guiding GFlowNet’s backward policy, which we utilized to improve computational efficiency by removing costly parent state computation, and to focus sampling on molecules with low synthesis cost, (2) Dynamic Library building blocks, which improve exploration efficiency—especially with smaller fragment libraries—and enable the generation of non-linear synthesis trees, and (3) an Exploitation Penalty, which avoids repeated visits to previously explored candidates. Through empirical evaluation, we demonstrated that combining these strategies leads to state-of-the-art results in template-based molecular generation across diverse experimental settings, from small-scale fragment libraries suitable for rapid, automated synthesis to large-scale libraries that create vast search spaces.
6 Limitations
While our study demonstrates the effectiveness of SCENT in generating cost-efficient and synthesizable molecules, several limitations remain.
First, the synthesis cost estimates are simplified and heuristic, omitting factors such as solvent usage, purification steps, hazardous reagents, reaction conditions, and scalability. Consequently, the current cost model should be regarded as a first-order approximation. More accurate estimation would require lab-specific calibration, including robust yield prediction models and vendor-specific building block pricing.
SCENT’s performance varies across different biological targets, indicating that its effectiveness may depend on task-specific factors. Identifying and addressing these bottlenecks is an important direction for future work. Key areas of improvement include refining the exploitation behavior of the Dynamic Library, and enhancing the generalizability of the cost predictor model used in Synthesis Cost Guidance. Additionally, we do not assess performance as the number of reaction templates increases, an important scalability consideration, as this expansion significantly enlarges the action space and introduces additional learning challenges.
Our framework emphasizes synthesizability and cost, but does not incorporate toxicity or broader ADMET properties, which are critical for downstream drug development. Finally, we acknowledge the absence of wet-lab validation, which remains the definitive benchmark for evaluating the practical utility of our approach. We leave this to future work.
Acknowledgements
The research of P. Gaiński was supported by by the National Science Centre (Poland), grant no. 2022/45/B/ST6/01117. O. Boussif was supported by the National Research Council Canada (NRC) AI4D program. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017716. The research was also enabled by computational resources provided by the Digital Research Alliance of Canada (https://alliancecan.ca) and Mila (https://mila.quebec). We thank Marcin Sendera and Moksh Jain for their support during the early stages of developing this paper.
@@ A.3.1 Synthesis Cost Model
We train the model $\hat{c}_{B}^{S}$ using the procedure described in Section˜3.2. Given an updated dataset with synthesis cost estimates $D$, we perform five mini-batch updates of $\hat{c}_{B}^{S}$ using mini-batches of size $1024$, the learning rate of $0.01$ and the mean square error loss $\mathcal{L}_{cost}=\sum_{c_{s}\in D^{\prime}}\lvert\lvert c_{s}-\hat{c}_{B}^{S}\rvert\rvert_{2}^{2}$, where $D^{\prime}$ is a mini-batch of $D$. To stabilize training, we standardize all costs in $D$ using the mean and variance of costs from the building block library $M$.
@@ A.3.2 Decomposability Model
The decomposability model $\hat{c}_{B}^{D}$ is very similar to the synthesis cost model, except that it encodes the current number of reactions used in the molecule $s$. The "cost" in the context of decomposability is 0 if the given molecule can be decomposed into building blocks from $M$ (within some number of steps) and $1$ otherwise. In practice, we train the model $\hat{c}_{B}^{D}$ to predict the decomposability label $\in\{0,1\}$ using binary cross entropy $\mathcal{L}_{cost}$. The labels are gathered similarly to the synthesis cost values, using the trajectories encountered in the GFlowNet forward sampling. We update $\hat{d}(s)$ with five minibatches of size $1024$ and a learning rate of $0.005$. In addition, we ensure that at least $20\%$ of the samples of molecules for training $\hat{d}(s)$ are not decomposable.
@@ B.1 SCENT
All SCENT instances were trained using Decomposability Guidance, either solely or with Synthesis Cost Guidance using Equation˜11.
@@
To predict the decomposability of the molecule, we use the model described in Section˜A.3. We set the cost temperature (Equation˜10) to $\alpha_{D}=5$.
@@
The model for Synthesis Cost Guidance is described in Section˜A.3. We set the cost temperature to $\alpha=5$.
@@
We set $\epsilon=3$ in Equation˜12 and schedule $\gamma$ in two ways: 1) so that it linearly decays to zero after $N$ iterations, and 2) it grows with the trajectory length (each step increases it by some constant factor $\Delta\gamma$). We set $N=\infty$ for SCENT that uses Dynamic Library and $N=1000$ for the rest of the runs. The growing delta $\Delta\gamma=0.2$, while the initial $s_{0}$ temperature $\gamma_{0}=1.0$. Note that for JNK3 in the MEDIUM setting, we used $4500$ iterations for exploitation penalty.
@@ B.4 RxnFlow
Appendix C Hyperparameters Sensitivity
To assess SCENT (C+D+P) sensitivity to hyperparameters, we performed an ablation study on sEH MEDIUM, varying each hyperparameter around the chosen values (center in tables) over 3k iterations. Results in Table˜2 indicate SCENT is generally robust, with no catastrophic performance drops. However, some hyperparameters notably affect outcomes; for example, simple tuning of $\alpha_{D}$ can boost discovered modes by  20%. This highlights potential for further improvements through targeted tuning, left for future work.
Table 2: Sensitivity of SCENT (C+D+P) to hyperparameters. Each hyperparameter was varied around the chosen values (center in tables). The results are reported for MEDIUM settings after 3k training iterations.
model online discovery inference sampling
modes >> 8.0 $\uparrow$ scaff. >> 8.0 $\uparrow$ reward $\uparrow$ cost >> 8.0 $\downarrow$
Synthesis Cost Guidance (temperature $\alpha_{S}$)
$\alpha_{S}=2$ 16036 ± 1153 22292 ± 1652 7.55 ± 0.05 1300 ± 82
$\alpha_{S}=5$ 16373 ± 1827 39721 ± 3574 7.61 ± 0.03 1163 ± 81
$\alpha_{S}=10$ 13566 ± 4207 73916 ± 2516 7.87 ± 0.04 898 ± 42
Decomposability Guidance (temperature $\alpha_{D}$)
$\alpha_{D}=2$ 20023 ± 1484 45483 ± 2490 7.74 ± 0.04 1137 ± 130
$\alpha_{D}=5$ 16373 ± 1827 39721 ± 3574 7.61 ± 0.03 1163 ± 81
$\alpha_{D}=10$ 20444 ± 2055 44938 ± 3111 7.74 ± 0.08 1141 ± 80
Dynamic Library (number of added molecules $L$)
$L=200$ 19700 ± 1919 49119 ± 6425 7.72 ± 0.05 1087 ± 183
$L=400$ 16373 ± 1827 39721 ± 3574 7.61 ± 0.03 1163 ± 81
$L=800$ 18373 ± 987 48016 ± 4302 7.71 ± 0.08 1060 ± 66
Dynamic Library (frequency of updates $T$)
$T=500$ 15670 ± 2482 39701 ± 3676 7.66 ± 0.12 1130 ± 106
$T=1000$ 16373 ± 1827 39721 ± 3574 7.61 ± 0.03 1163 ± 81
$T=2000$ 16148 ± 1753 39099 ± 3341 7.66 ± 0.08 1141 ± 82
Exploitation Penalty (initial $\gamma_{0}$)
$\gamma_{0}=0.5$ 20690 ± 1512 47220 ± 3676 7.69 ± 0.06 1154 ± 48
$\gamma_{0}=1.0$ 16373 ± 1827 39721 ± 3574 7.61 ± 0.03 1163 ± 81
$\gamma_{0}=2.0$ 20655 ± 1514 45990 ± 1593 7.75 ± 0.04 1111 ± 77
Exploitation Penalty ($\Delta\gamma$)
$\Delta\gamma=0.0$ 20025 ± 762 43432 ± 796 7.75 ± 0.02 1222 ± 65
$\Delta\gamma=0.2$ 16373 ± 1827 39721 ± 3574 7.61 ± 0.03 1163 ± 81
$\Delta\gamma=0.4$ 17189 ± 2234 42718 ± 2799 7.69 ± 0.04 1108 ± 165
Appendix D Fragment analysis
In this section, we showcase characteristics of the fragments we used in this paper for all settings. We focus on molecular weight, lipophilicity (logP), the number of Hydrogen bond donors (HBD) and acceptors (HBA), the Topological polar surface area (TPSA) and the number of rotatable bonds (See Figure 8).
Figure 8: Exploratory analysis of chemical properties of fragments in the SMALL, MEDIUM and LARGE settings.
Appendix E Reaction Yields and Fragments Costs Estimates
@@
Building block prices for SMALL were obtained from the vendors online. If multiple vendors offered the same building block, only the cost of the cheapest option was considered. For compounds available in varying amounts, both the price per gram and the alternative size (e.g., 5, 10, 25 grams) were considered, and the smallest price per mmol was used. Compounds that were exclusively accessible in other forms, such as salts, related functional groups, or containing a protected functional group were used with the SMILES fitted to the corresponding reactions. The cost of a product includes the total cost of the building blocks used, while additional expenses such as solvents, catalysts, and reagents are not accounted for. For the MEDIUM and LARGE settings, stock prices were obtained by automatically scraping Enamine’s publicly available catalog.
For reactions where sufficient in-house experimental data was available (amide coupling, nucleophilic additions to isocyanates, Suzuki reaction, Buchwald–Hartwig coupling, Sonogashira coupling, and azide–alkyne cycloaddition), yields were calculated as the average of all recorded experimental yields corresponding to each SMARTS-based reaction template. For the remaining reactions in the dataset, yield estimates were derived from the average of reported literature yields for reactions matching the respective SMARTS templates.
@@
Appendix F Extended Comparison to Template-Based GFlowNets
F.1 Results on sEH with Reward Threshold = 7.2
Despite tuning the SynFlowNet [9] parameters (Appendix˜B), we could not obtain satisfying results. Importantly, our experimental setting differs from the one used in their paper in the following way: 1) we use up to 4 reactions, they use 2 or 3, 2) they use a 10k subset of Enamine, our MEDIUM setting contains 64k molecules, 3) we set the reward thresholds higher (for sEH we set 8.0 instead of 7.2), and 4) we train the model for fewer trajectories. In Table˜3, we report the results for sEH using a threshold from the SynFlowNet paper. The number of modes obtained seems consistent with those reported in Cretu et al. [9].
Table 3: Evaluation of different template-based approaches using sEH proxy in SMALL, MEDIUM and LARGE settings, but using lower reward threshold (>7.2>7.2) equivalent to the one used in SynFlowNet paper [9].
@@ Appendix F Extended Comparison to Template-Based GFlowNets
modes >> 7.2 $\uparrow$ scaff. >> 7.2 $\uparrow$ reward $\uparrow$ cost >> 7.2 $\downarrow$
SMALL
RGFN 11320 $\pm$168 76742 $\pm$693 7.34 $\pm$0.06 35.3 $\pm$1.6
@@
modes >> 0.8 $\uparrow$ scaff. >> 0.8 $\uparrow$ reward $\uparrow$ cost >> 0.8 $\downarrow$
SMALL
RGFN 285 $\pm$14 1870 $\pm$106 0.7 $\pm$0.01 42.3 $\pm$2.7
@@
Table 5: Evaluation of different template-based approaches using JNK3 proxy in SMALL, MEDIUM and LARGE settings. In all metrics, except "reward", we consider molecules with reward >0.8>0.8. We observe that SCENT (C) that uses Cost Guidance can greatly reduce the cost of the sampled molecules in SMALL and MEDIUM settings, while drastically increasing the number of scaffolds.
@@
modes >> 0.8 $\uparrow$ scaff. >> 0.8 $\uparrow$ reward $\uparrow$ cost >> 0.8 $\downarrow$
SMALL
RGFN 307 $\pm$11 2852 $\pm$42 0.8 $\pm$0.0 11.6 $\pm$0.3
@@ F.3 Results on JNK3
In Table˜5 RGFN outperforms all SCENT variants in the smallest setting, consistently finding more diverse modes and high-reward scaffolds. While SCENT lags behind in terms of mode discovery, it compensates in terms of cost, since SCENT (C) samples the cheapest molecules. In the MEDIUM setting, SCENT (w/o C) outperforms the baseline methods and the fully enabled SCENT variants in terms of mode discovery, but lags behind in the number of high-reward scaffolds. That can be attributed to the fact that SCENT (C) samples are less diverse. However, SCENT (C+P) samples the cheapest molecules. Finally, in the largest setting, the fully enabled SCENT (C+P) finds the most number of modes and high-reward scaffolds, whereas SCENT (C) takes the spotlight in terms of cost, generating the cheapest molecules.
F.4 Comparison of computational resources
In Table˜6 report GPU runtimes (in minutes) for all template-based baselines on the sEH proxy, averaged over 3 random seeds using a V100 32GB GPU. All methods used a batch size of 64, except SynFlowNet, which required batch size 32 on the LARGE setting even after optimizing its policy code to fit within memory limits.
RGFN shares a part of the codebase with SCENT, allowing for direct comparison. We observe consistent improvements across all settings, particularly in the SMALL configuration. While SynFlowNet is faster on SMALL and MEDIUM, it struggles on LARGE, even with our memory optimizations. However, differences in codebases make direct comparisons less definitive.
Note that our implementation is not optimized for runtime and could benefit from further improvements.
Table 6: GPU runtimes (in minutes) for all template-based baselines on the sEH proxy, averaged over 3 random seeds using a V100 32GB GPU.
Model SMALL MEDIUM LARGE
RGFN 2882 ± 115 3892 ± 153 4455 ± 392
SynFlowNet 200 ± 3 2350 ± 162 8416 ± 19
RxnFlow 891 ± 4 3304 ± 44 3399 ± 50
SCENT (C+D+P) 800 ± 15 2849 ± 144 3688 ± 317
Appendix G Qualitative Pathways Analysis
To visualize the Synthesis Cost Guidance in practice, we gathered training trajectories from SCENT (C) and SCENT (w/o C) that lead to the same high-rewarded molecule in the SMALL setting.
We observed that, on average, cost-guided trajectories are  10% cheaper. We selected a few example trajectory pairs and noticed that the cost reduction usually occurs by changing the order of the reactions: introducing expensive fragments later in the synthesis path, which decreases the product losses due to imperfect yields, and as a result increases the cost-efficiency. In Figure˜9, SCENT (C) chooses the most expensive building block in the sequence (\$12.56) at the beginning of the synthesis, while SCENT (C) utilizes it in the second step, effectively reducing the synthesis cost by almost 28%. Cost-guided model also occasionally prefers cheaper fragments, e.g., "Brc1c[nH]cn1" instead of "Ic1c[nH]cn1". It is worth emphasizing that the fragments in the SMALL setting are already low cost, which is the reason why most gains are observed due to reaction order swapping.
Similarly, we compared the Dynamic Library, SCENT (C+D), to SCENT (C), and observed that on average the trajectories are 12% shorter and 6% cheaper. We selected few examples that visualized the compressed trajectories. We have found that the Dynamic Library enables the generation of convergent synthetic pathways (Figure˜10), which are responsible for the reduction of the synthesis cost. The application of the convergent synthetic strategy decreases the longest linear sequence (LLS) of steps, thereby increasing the overall yield.
In the MEDIUM settings, the SCENT variations diverged early on, so the set of shared molecules is very limited. Therefore, we matched high-rewarded molecules from SCENT (C) to closest high-rewarded molecule from SCENT (w/o C) with Tanimoto similarity > 0.6. We observed that, on average, the cost-guided trajectories are 20% cheaper. Thorough investigation of few selected trajectories pairs revealed that, differently from SMALL setting, in MEDIUM setting, the cost reduction comes mostly from the selection of cheaper fragments (Figure˜11).
(a) SCENT w/o C
(b) SCENT C
(c) SCENT C+D
Figure 9: Three example pathways from SMALL setting leading to the same molecule. SCENT (C) is cheaper than SCENT (w/o C) by introducing the costly fragment (\$12.56) later in the pathway.
(a) SCENT w/o C
(b) SCENT C
(c) SCENT C+D
Figure 10: Three example pathways from SMALL setting leading to the same molecule. SCENT (C) is cheaper than SCENT (w/o C) by introducing the costly fragment (\$16.81) later in the pathway. SCENT (C+D) further lowers cost via a convergent strategy that shortens the longest linear sequence (LLS), boosting overall yield.
(a) SCENT w/o C
(b) SCENT C
Figure 11: Two example pathways leading to structurally similar high-rewarded molecules. A molecule proposed by SCENT (C) uses much cheaper fragment (\$27.27 vs \$279.35).
Appendix H Extended Synthesis Cost Guidance Analysis